# MS-0000-0000.00

### v1

@@ 4.1.3   Mechanisms and Resolutions
Shaping Human Expectations:  Human expectations for AI as a team member will likely be influenced through deployment and feedback mechanisms, which themselves be interdependent. Feedback mechanisms can be broadly categorized into (1) Descriptive feedback: AI provides neutral, observational feedback providing information without judgment (e.g., summarizing data trends). Feedback supports situational awareness but may not directly guide decision-making. (2) Normative feedback: AI makes goal-oriented suggestions emphasizing desired outcomes (e.g., recommending optimal actions). This feedback drives alignment with strategic objectives but may risk overdependence if the feedback is overly prescriptive. (3) Corrective feedback: Error-oriented adjustments highlighting execution and reasoning inaccuracies or deviations (e.g., pointing out flaws in task assignments and execution). While corrective feedback enhances accuracy, it requires careful framing to avoid undermining human confidence. (4) Evaluative feedback: Judgment-based assessments offering qualitative or quantitative evaluations (e.g., performance ratings). This feedback fosters accountability but can introduce biases if not transparently derived and delivered with a keen awareness of context. Humans absorb these feedback types through the integration of multiple channels (e.g., visual, auditory, sensory). Deployment also determines where the AI team member functions on the continuum of operant and operand resources. As an operand, AI functions as a passive entity requiring human activation. Whereas, as an operant, AI functions as an independent entity that makes independent judgments and takes actions on its own. The task interdependencies between AI and the human agent can become very fluid when human-AI interactions dynamically adjust along this continuum in reaction to situational awareness. We currently have no theoretical grounding to suggest optimal configurations for deployment and feedback mechanisms, let alone the interactions between the two or the potential possibility for dynamism in the mechanisms during task execution. For example, future studies can explore (1) constraints and ethical considerations related to feedback and deployment configurations, (2) impact on human learning and adaptability under different configurations, and (3) human expectations and experiences related to cognitive fluency and burden associated with AI agents moving from being an operand resource to an operant resource.
@@
Processing Transformative AI Outputs:  It is essential to understand how humans perceive, process, and learn from transformative AI outputs (e.g., suggestions, conclusions). An advanced AI teammate could break information bubbles and introduce novel perspectives for humans in their teamwork (Nguyen et al. 2014). This process involves analyzing how humans interpret AI-generated insights and the subsequent impact on their problem-solving approaches. There is a growing body of research that shows humans preferring AI for more objective decision contexts and preferring recommendations from another human for more subjective/complex decisions (Steyvers and Kumar 2024). In this vein, (1) to what extent such preferences are a function of current AI capabilities and human perceptions as opposed to longer-term behavioral preferences will remain an open question for the next few years. Transparency and explainability in recommendations can greatly influence the acceptance and utilization of AI recommendations in team processes. Furthermore, (2) to what extent humans believe in the neutrality of AI agents (i.e., not biased through managerial manipulations or organizational priorities that are not in alignment with human perception) can also be influential in how humans accept transformative inputs from AI agents. In addition to cognitive and perceptual outcomes, (3) studies that record interaction data to unpack the deliberation processes between humans and intelligent agents remain underexplored. This is particularly critical in the era of generative AI, as the performance of these agents heavily depends on how individuals utilize them and interpret the recommendations within the larger socio-technical context of their work.
@@ 4.2.1   Expected Outcomes
The integration of AI agents with SA capabilities holds great promise for enhancing team collaboration. By interpreting instructions, understanding objectives, and anticipating needs, assigning specialized roles to AI and integrating it into appropriate workflows can enhance task efficiency, optimize resource allocation, and improve decision-making. The growing autonomy of AI agents, however, introduces challenges. When crafting AI with strong learning capabilities to understand tasks, process feedback, and adapt to dynamic contexts, it is crucial to recognize key limitations. AI specification problem in a team context spans several issues. Most AI systems (even generative AI) perform well in data-rich environments. While some contexts are naturally data-rich, in other contexts, deliberate design for data collection and aggregation is a must for developing reliable AI agents. In socially complex and subjective decision-making scenarios, utilization of AI agents requires striking a balance between the tacit expertise of humans and rational insights from the AI agents (Sharma et al. 2023).
@@ 4.2.2   Evaluations and Issues
The AI Role Specification Challenge:  The understanding of team activities by AI can be broadly viewed as a challenge of role specification. The clarity and precision of the specifications governing both product and process characteristics in the team environment determine how well human-to-AI knowledge transfer is facilitated. Product characteristics encompass the attributes, requirements, and goals of the tasks or outputs the AI is expected to deliver, while process characteristics define workflows, rules, and interaction protocols that structure team collaboration with AI. Robust specifications are essential for ensuring that AI systems accurately perceive (Level 1) and comprehend (Level 2) their environment. When specification is imperfect, AI may struggle with accuracy, misinterpret ambiguous human inputs, or fail to adapt to changing objectives. Defining the parameters of AI role specification within the team is essential to enhancing its SA perception and the resulting interaction within the team. Future studies should investigate (1) how well AI systems identify task requirements and interpret human prompts, particularly in ambiguous scenarios involving both intentional and unintentional inputs. This includes examining the factors that influence AI’s accuracy in recognizing and adapting to dynamic team configurations. Additionally, (2) researchers should explore how AI systems process evolving objectives and integrate these changes into workflows to maintain alignment with team goals. These efforts will inform the development of more adaptive role specifications, ensuring that AI systems can respond effectively to the complexity and fluidity of human-AI collaboration (Teaming 2022). It is also possible to envision AI itself identifying role ambiguity when deployed and may even guide the team to reconfigure team processes and specifications.
@@ 4.3 Social Specification: Human–AI Team Formulation and Coordination
A human-AI team operates as a socio-technical system, requiring well-crafted social specifications to achieve cohesion and efficacy. Specifications such as defined roles, rules, and norms are needed to guide team formulation, coordination, and interaction. Particularly, as discussed before, human-AI teaming introduces new dynamics that demand modifications to engagement modes to integrate AI agents to achieve collective goals. Organizations that excel in managing these specifications will outperform their counterparts, leveraging both human and AI strengths effectively. Below, we explore the expected outcomes, key evaluations and issues, and mechanisms for addressing opportunities and challenges in this domain, while proposing specific research directions for future exploration.
@@ 4.3.1   Expected Outcomes
Human-AI teams can achieve synergy by combining AI’s vast computational power with human intuition and expertise. This integration can enhance task execution, adaptability, and innovation. An expected operational outcome from optimized role specification allows team members, both human and AI, to leverage their strengths, resulting in faster decision-making, higher-quality outcomes, and greater satisfaction (Panico 2024). It is in more complex and creative team tasks there are new opportunities for researchers to explore. While the focus has been on easing cognitive burden and augmenting cognitive capabilities, new research on social specifications that foster cognitive leaps is still scarce. We assert that cognitive leaps at current AI capabilities are only possible with humans in the loop. But AI support can greatly aid in humans achieving cognitive leaps in their tasks. By definition, cognitive leaps are discontinuous reasoning or insights as opposed to stepwise logical or linear reasoning. Helping humans connect concepts and data that at the surface seem unrelated can greatly aid cognitive leaps. Moreover, when the cognitive burden on routine reasoning tasks is eased, it enables humans to focus on more complex concepts and tasks. Social specification for human-AI teaming can hinder or encourage cognitive leaps through strategies that identify deeply hidden patterns in data, exploration of adjacent ideas and domains, synthesizing disjoint knowledge streams, and others. Achieving synergies is as much a technical problem (e.g., interaction design, data identification, model support) as it is a social problem (e.g., recognizing and rewarding human-AI innovations, autonomy for decision-making, ability to incorporate outside data sources). This includes balancing accountability between human and AI members and developing mechanisms to value diverse contributions effectively (Teaming 2022, Grote et al. 2024).
@@ 4.3.2   Evaluations and Issues
Task Work, Role, and Resource Allocation Effectiveness:  Task and role allocation are critical for ensuring human-AI team synergy. Synergistic allocations effectively blend AI’s computational strengths and human insights to achieve team objectives. The reason we identify social specification issues here is that for humans to effectively challenge/alter/recombine AI recommendations, they should be enabled to do so. That is, human team members who override or refine AI recommendations should be enabled and encouraged to take such actions. Yet, there should be a mechanism to evaluate their decisions against actual outcomes. The feedback mechanisms, however, need careful calibration lest they create risk-aversive behavior in the team. Future research should (1) explore methods for optimal team designs that encourage humans to contextualize AI recommendations, enable AI agents to discover deep patterns, and provide feedback mechanisms that improve both human and AI performance over time. In more creative and explorative task contexts, future research can (2) identify how AI can accelerate choice generation (such as prototypes and policy frameworks), and humans can weigh in to filter and refine the choice sets based on organizational context and competitive strategies. Additionally, researchers should (3) examine the scalability of role allocation strategies in larger and more diverse human-AI teams. For instance, as team size and complexity increase, new challenges may arise in ensuring that roles remain clear, well-defined, and adaptable. Studying how task allocation systems perform across varying team sizes and contexts can help develop robust frameworks that maintain efficiency and clarity in increasingly complex environments.
@@
Responsibility and Rewards:   Effective evaluation of responsibility assignments and reward systems is essential for understanding team motivation, fairness, and the successful integration of role specification in human-AI collaboration. The implications of human-AI teaming on reward and performance systems will be an important research focus for a number of reasons. First, reward systems have tended to recognize individual contributions, productivity, and effort. In most cases, such reward systems tend to work on the assumption that skills and effort are the primary drivers of performance. Human-AI teaming fundamentally alters this calculus where intuition, critical thinking, and creativity will have a large impact on performance (with AI providing the skills input in most cases). Second, the infusion of AI agents in teams will alter the reporting structure considerably. Predictions of the need for fewer managers in an AI-augmented world are almost uniformly acknowledged. At the minimum, this would lead to managers having a wider scope of oversight over multiple team members. The ability of managers to be able to recognize creative thinking among the human team members can be severely curtailed in such a context. Clearly, new performance assessment mechanisms will be needed to equitably assess team members. Clearly defined responsibilities that leverage the unique strengths of both human and AI members enhance accountability and efficiency by ensuring each member contributes according to their specification. Conversely, poorly defined roles can lead to ambiguity, misaligned expectations, and inefficiencies, while imbalanced reward systems risk demotivating team members and undermining cohesion (Teaming 2022). Third, any changes to performance measurement systems will instill feedback mechanisms that, in turn, dictate how human-AI teams perform. Therefore, assessing the efficacy of any performance measurement systems will be hard in the near term for any team. Future studies should (1) explore innovative methods to balance responsibility and rewards in human-AI teams, integrating role specification to ensure accountability and equitable recognition of contributions. (2) Mechanisms such as decentralized governance models and stakeholder negotiation frameworks (Grote et al. 2024) offer promising directions to enhance fairness and foster cohesion, ultimately optimizing team performance.
@@ 4.3.3   Mechanisms and Resolutions
Optimizing Role Specification and Adaptability:  Role specification ensures that team members, both human and AI, can leverage their strengths to contribute effectively to team goals. In human-AI teaming, AI’s computational capabilities must complement human intuition and contextual understanding. However, rigid role structures may fail to adapt to dynamic task demands, necessitating more flexible approaches to role assignment (Marks et al. 2000). Adaptive algorithms that adjust roles based on real-time performance and task requirements can enhance team effectiveness (Zhang et al. 2024a, Raisch and Fomina 2023). Future research should (1) explore the development of frameworks for determining when AI should take on specific responsibilities, such as assisting in intermediate stages or making final decisions (Yin et al. 2025). Additionally, researchers should (2) investigate how collaborative modes, such as parallel versus sequential workflows, impact team dynamics and outcomes (Chen and Chan 2024, Choudhary et al. 2025).
@@
Adopting A Socio-Technical Perspective on Performance Evaluation:  The socio-technical perspective provides a holistic framework for evaluating team performance by examining AI’s impacts and key decisions across its lifecycle. It emphasizes assessing AI’s suitability for tasks while addressing the needs of individuals, groups, and society (Schwartz et al. 2022). By analyzing the interplay between human and technological factors, such as social dynamics, individual skills, tools, and organizational contexts, this perspective highlights drivers and barriers to team effectiveness. Addressing challenges like AI-induced job insecurity and maladaptive workplace behaviors is essential for fostering resilience and collaboration in human-AI teams (Yam et al. 2023). Future studies should thus (1) investigate AI’s socio-psychological impacts on trust and motivation to uncover pathways for improving team performance and cohesion. Moreover, fostering effective human-AI collaboration also requires addressing job security, emotional support, and adaptive teamwork. The perceived threat of AI replacing jobs undermines team cohesion and warrants urgent attention (Yam et al. 2023). Game-theoretical models can provide insights into balancing human-AI roles for equitable outcomes (Panico 2024). Skill-enhancing training and equitable work distribution are critical strategies for mitigating these risks. Similarly, experienced workers’ concerns about maintaining their unique contributions must be addressed, as advanced AI may reduce expertise gaps, potentially disadvantaging highly skilled individuals (Brynjolfsson and Mitchell 2017, SimanTov-Nachlieli 2024). Future studies should (2) develop AI systems with adaptive communication skills, such as interpreting social signals and recognizing personality traits (Li and Bitterly 2024, Gabriel et al. 2024), to enhance trust and collaboration in human-AI teams.
@@
Designing Effective Incentive and Credit Structures:  Well-designed incentive and credit systems are vital for maintaining motivation and fairness in human-AI teams. Poorly designed structures can lead to disengagement, reduced morale, and diminished collaboration. Performance-based incentives, recognition programs, and credit systems that acknowledge contributions from both human and AI members are critical for fostering active participation and accountability (Teaming 2022). Additionally, mechanisms to align individual goals with team objectives can promote collective responsibility and cohesion (Geng and Varshney 2022). Future research should (1) explore innovative incentive models that balance financial and non-financial rewards, enhancing commitment and motivation in human-AI teams. Future studies should also (2) explore the role of perceived equity in the distribution of responsibilities (perhaps more inline with creativity and insights needed in the tasks) and rewards, examining how differences in team members’ roles and contributions – especially between humans and AI – affect perceptions of fairness and team cohesion. For example, investigating whether teams with highly specialized AI roles require distinct reward and accountability structures compared to those with more balanced human-AI contributions could offer new insights into designing fair and effective evaluation systems.

### v2

@@ 4.1.3   Mechanisms and Resolutions
Shaping Human Expectations:  Human expectations for AI as a team member will likely be influenced through deployment and feedback mechanisms, which themselves be interdependent. Feedback mechanisms can be broadly categorized into (1) Descriptive feedback: AI provides neutral, observational feedback providing information without judgment (e.g., summarizing data trends). Feedback supports situational awareness but may not directly guide decision-making. (2) Normative feedback: AI makes goal-oriented suggestions emphasizing desired outcomes (e.g., recommending optimal actions). This feedback drives alignment with strategic objectives but may risk overdependence if the feedback is overly prescriptive. (3) Corrective feedback: Error-oriented adjustments highlighting execution and reasoning inaccuracies or deviations (e.g., pointing out flaws in task assignments and execution). While corrective feedback enhances accuracy, it requires careful framing to avoid undermining human confidence. (4) Evaluative feedback: Judgment-based assessments offering qualitative or quantitative evaluations (e.g., performance ratings). This feedback fosters accountability but can introduce biases if not transparently derived and delivered with a keen awareness of context. Humans absorb these feedback types through the integration of multiple channels (e.g., visual, auditory, sensory). Deployment also determines where the AI team member functions on the continuum of operant and operand resources. As an operand, AI functions as a passive entity requiring human activation. Whereas, as an operant, AI functions as an independent entity that makes independent judgments and takes actions on its own. The task interdependencies between AI agents and humans can become very fluid when human-AI interactions dynamically adjust along this continuum in reaction to situational awareness. We currently have no theoretical grounding to suggest optimal configurations for deployment and feedback mechanisms, let alone the interactions between the two or the potential possibility for dynamism in the mechanisms during task execution. For example, future studies can explore (1) constraints and ethical considerations related to feedback and deployment configurations, (2) impact on human learning and adaptability under different configurations, and (3) human expectations and experiences related to cognitive fluency and burden associated with AI agents moving from being an operand resource to an operant resource.
@@
Processing Transformative AI Outputs:  It is essential to understand how humans perceive, process, and learn from transformative AI outputs (e.g., suggestions, conclusions). An advanced AI teammate could break information bubbles and introduce novel perspectives for humans in their teamwork (Nguyen et al. 2014). This process involves analyzing how humans interpret AI-generated insights and the subsequent impact on their problem-solving approaches. There is a growing body of research that shows humans preferring AI for more objective decision contexts and preferring recommendations from another human for more subjective/complex decisions (Steyvers and Kumar 2024). In this vein, (1) to what extent such preferences are a function of current AI capabilities and human perceptions as opposed to longer-term behavioral preferences will remain an open question for the next few years. Transparency and explainability in recommendations can greatly influence the acceptance and utilization of AI recommendations in team processes. Furthermore, (2) to what extent humans believe in the neutrality of AI agents (i.e., not biased through managerial manipulations or organizational priorities that are not in alignment with human perception) can also be influential in how humans accept transformative inputs from AI agents. In addition to cognitive and perceptual outcomes, (3) studies that record interaction data to unpack the deliberation processes between humans and AI agents remain underexplored. This is particularly critical in the era of generative AI, as the performance of these agents heavily depends on how individuals utilize them and interpret the recommendations within the larger socio-technical context of their work.
@@ 4.2.1   Expected Outcomes
The integration of AI agents with SA capabilities holds great promise for enhancing team collaboration. By interpreting instructions, understanding objectives, and anticipating needs, assigning specialized roles to AI and integrating it into appropriate workflows can enhance task efficiency, optimize resource allocation, and improve decision-making. The growing autonomy of AI agents, however, introduces challenges. When crafting AI with strong learning capabilities to understand tasks, process feedback, and adapt to dynamic contexts, it is crucial to recognize key limitations. AI specification problem in a team context spans several issues. Modern AI systems perform well in data-rich environments. While some contexts are naturally data-rich, in other contexts, deliberate design for data collection and aggregation is a must for developing reliable AI agents. In socially complex and subjective decision-making scenarios, utilization of AI agents requires striking a balance between the tacit expertise of humans and rational insights from the AI agents (Sharma et al. 2023).
@@ 4.2.2   Evaluations and Issues
The AI Role Specification Challenge:  The understanding of team activities by AI can be broadly viewed as a challenge of role specification. The clarity and precision of the specifications governing both product and process characteristics in the team environment determine how well human-to-AI knowledge transfer is facilitated. Product characteristics encompass the attributes, requirements, and goals of the tasks or outputs the AI is expected to deliver, while process characteristics define workflows, rules, and interaction protocols that structure team collaboration with AI. Robust specifications are essential for ensuring that AI systems accurately perceive (Level 1) and comprehend (Level 2) their environment. When specification is imperfect, AI may struggle with accuracy, misinterpret ambiguous human inputs, or fail to adapt to changing objectives. Defining the parameters of AI role specification within the team is essential to enhancing its SA perception and the resulting interaction within the team. Future studies should investigate (1) how well AI systems identify task requirements and interpret human prompts, particularly in ambiguous scenarios involving both intentional and unintentional inputs. This includes examining the factors that influence AI’s accuracy in recognizing and adapting to dynamic team configurations. Additionally, (2) researchers should explore how AI systems process evolving objectives and integrate these changes into workflows to maintain alignment with team goals. These efforts will inform the development of more adaptive role specifications, ensuring that AI systems can respond effectively to the complexity and fluidity of human-AI collaboration (Teaming 2022).
@@ 4.3 Social Specification: Human–AI Team Formulation and Coordination
A human-AI team operates as a socio-technical system, requiring well-crafted social specifications to achieve cohesion and efficacy. Specifications such as defined roles, rules, and norms are needed to guide team formulation, coordination, and interaction. Particularly, as discussed before, human-AI teaming introduces new dynamics that demand modifications to engagement modes to integrate AI agents to achieve collective goals. Organizations that excel in managing these specifications will outperform their counterparts, leveraging both human and AI strengths effectively. Below, as summarized in Figure 3, we explore the expected outcomes, key evaluations and issues, and mechanisms for addressing opportunities and challenges in this domain, while proposing specific research directions for future exploration.
[…]
Figure 3: A Framework for Team Formulation and Coordination in Human–AI Teams
[…]
@@ 4.3.1   Expected Outcomes
Human-AI teams can achieve synergy by combining AI’s vast computational power with human intuition and expertise. This integration can enhance task execution, adaptability, and innovation. Optimized role specification allows team members, both human and AI, to leverage their strengths, resulting in faster decision-making, higher-quality outcomes, and greater satisfaction (Panico 2024). Greater opportunities for research emerge in more complex and creative team tasks. While the focus has been on easing cognitive burden and augmenting cognitive capabilities, new research on social specifications that foster cognitive leaps is still scarce. We assert that cognitive leaps at current AI capabilities are only possible with humans in the loop (Choudhary et al. 2025). But AI support can greatly aid in humans achieving cognitive leaps in their tasks. By definition, cognitive leaps are discontinuous reasoning or insights as opposed to stepwise logical or linear reasoning. Helping humans connect concepts and data that at the surface seem unrelated can greatly aid cognitive leaps. Moreover, when the cognitive burden on routine reasoning tasks is eased, it enables humans to focus on more complex concepts and tasks. Social specification for human-AI teaming can hinder or encourage cognitive leaps through strategies that identify deeply hidden patterns in data, exploration of adjacent ideas and domains, synthesizing disjoint knowledge streams, and others. Achieving synergies is as much a technical problem (e.g., interaction design, data identification, model support) as it is a social problem (e.g., recognizing and rewarding human-AI innovations, autonomy for decision-making, ability to incorporate outside data sources). This includes balancing accountability between human and AI members and developing mechanisms to value diverse contributions effectively (Teaming 2022, Grote et al. 2024).
@@ 4.3.2   Evaluations and Issues
Task Work, Role, and Resource Allocation Effectiveness:  Task and role allocation are critical for ensuring human-AI team synergy. Synergistic allocations effectively blend AI’s computational strengths and human insights to achieve team objectives. The reason we identify social specification issues here is that for humans to effectively challenge/alter/recombine AI recommendations, they should be enabled to do so. That is, human team members who override or refine AI recommendations should be enabled and encouraged to take such actions. Yet, there should be a mechanism to evaluate their decisions against actual outcomes. The feedback mechanisms, however, need careful calibration lest they create risk-aversive behavior in the team. Future research should (1) explore methods for optimal team designs that encourage humans to contextualize AI recommendations, enable AI agents to discover deep patterns, and provide feedback mechanisms that improve both human and AI performance over time. In more creative and explorative task contexts, future research can (2) identify how AI can accelerate choice generation (such as prototypes and policy frameworks), and humans can weigh in to filter and refine the choice sets based on organizational context and competitive strategies.
@@
Responsibility and Rewards:   Effective evaluation of responsibility assignments and reward systems is essential for understanding team motivation, fairness, and the successful integration of role specification in human-AI collaboration. The implications of human-AI teaming on reward and performance systems will be an important research focus for a number of reasons. First, reward systems have tended to recognize individual contributions, productivity, and effort. In most cases, such reward systems tend to work on the assumption that skills and effort are the primary drivers of performance. Human-AI teaming fundamentally alters this calculus where intuition, critical thinking, and creativity will have a large impact on performance (with AI providing the skills input in most cases). Second, the infusion of AI agents in teams will alter the reporting structure considerably. Predictions of the need for fewer managers in an AI-augmented world are almost uniformly acknowledged. At the minimum, this would lead to managers having a wider scope of oversight over multiple team members. The ability of managers to be able to recognize creative thinking among the human team members can be severely curtailed in such a context. Clearly, new performance assessment mechanisms will be needed to equitably assess team members. Clearly defined responsibilities that leverage the unique strengths of both human and AI members enhance accountability and efficiency by ensuring each member contributes according to their specifications. Conversely, poorly defined roles can lead to ambiguity, misaligned expectations, and inefficiencies, while imbalanced reward systems risk demotivating team members and undermining cohesion (Teaming 2022). Third, any changes to performance measurement systems will instill feedback mechanisms that, in turn, dictate how human-AI teams perform. Therefore, assessing the efficacy of any performance measurement systems will be hard in the near term for any team. Future studies should (1) explore innovative methods to balance responsibility and rewards in human-AI teams, integrating role specification to ensure accountability and equitable recognition of contributions. (2) Mechanisms such as decentralized governance models and stakeholder negotiation frameworks (Grote et al. 2024) offer promising directions to enhance fairness and foster cohesion, ultimately optimizing team performance.
@@ 4.3.3   Mechanisms and Resolutions
Optimizing Role Specification and Adaptability:  Role specification ensures that team members, both human and AI, can leverage their strengths to contribute effectively to team goals. In human-AI teaming, AI’s computational capabilities must complement human intuition and contextual understanding. However, rigid role structures may fail to adapt to dynamic task demands, necessitating more flexible approaches to role assignment (Marks et al. 2000). Adaptive algorithms that adjust roles based on real-time performance and task requirements can enhance team effectiveness (Zhang et al. 2024a, Raisch and Fomina 2023). We elaborate further on role fluidity in Section 4.4. Future research should (1) explore the development of frameworks for determining when AI should take on specific responsibilities, such as assisting in intermediate stages or making final decisions (Yin et al. 2025). Additionally, researchers should (2) investigate how collaborative modes, such as parallel versus sequential workflows, impact team dynamics and outcomes (Chen and Chan 2024, Choudhary et al. 2025).
@@
Adopting A Socio-Technical Perspective on Performance Evaluation:  The socio-technical perspective provides a holistic framework for evaluating team performance by examining AI’s impacts and key decisions across its lifecycle. It emphasizes assessing AI’s suitability for tasks while addressing the needs of individuals, groups, and society (Schwartz et al. 2022). By analyzing the interplay between human and technological factors, such as social dynamics, individual skills, tools, and organizational contexts, this perspective highlights drivers and barriers to team effectiveness. Addressing challenges like AI-induced job insecurity and maladaptive workplace behaviors is essential for fostering resilience and collaboration in human-AI teams (Yam et al. 2023). Future studies should thus (1) investigate AI’s socio-psychological impacts on trust and motivation to improve team performance and cohesion. Moreover, fostering effective human-AI collaboration also requires addressing job security, emotional support, and adaptive teamwork. The perceived threat of AI replacing jobs undermines team cohesion and warrants urgent attention (Yam et al. 2023). Game-theoretical models could be developed to provide insights into balancing human-AI roles for equitable outcomes (Panico 2024). Skill-enhancing training and equitable work distribution are critical strategies for mitigating these risks. Similarly, experienced workers’ concerns about maintaining their unique contributions must be addressed, as advanced AI may reduce expertise gaps, potentially disadvantaging highly skilled individuals (Brynjolfsson and Mitchell 2017, SimanTov-Nachlieli 2024). Future studies should (2) develop AI systems with adaptive communication skills, such as interpreting social signals and recognizing personality traits (Li and Bitterly 2024, Gabriel et al. 2024), to enhance trust and collaboration in human-AI teams.
@@
Designing Effective Incentive and Credit Structures:  Well-designed incentive and credit systems are vital for maintaining motivation and fairness in human-AI teams. Poorly designed structures can lead to disengagement, reduced morale, and diminished collaboration. Performance-based incentives, recognition programs, and credit systems that acknowledge contributions from both human and AI members are critical for fostering active participation and accountability (Teaming 2022). Additionally, mechanisms to align individual goals with team objectives can promote collective responsibility and cohesion (Geng and Varshney 2022). Future research should explore (1) innovative incentive models that balance financial and non-financial rewards, enhancing commitment and motivation in human-AI teams, as well as (2) the role of perceived equity in the distribution of responsibilities (perhaps more inline with creativity and insights needed in the tasks) and rewards, delving into how differences in team members’ roles and contributions – especially between humans and AI – affect perceptions of fairness and team cohesion.